# TwinMarket: A Scalable Behavioral and Social Simulation for Financial Markets

### v2

@@ 4.2 Micro-analysis: Self-fulfilling Prophecy
Self-fulfilling prophecy refers to the phenomenon where market participants act based on a certain expectation, ultimately leading to the realization of that expectation [30].
In our BDI framework, we model users’ evolving expectations as belief, which is updated daily. The belief is structured along five key dimensions: the users’ view on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment (detailed in Appendix D.3). Each of these dimensions is assigned an emotional score (higher score means more positive) using GPT-4o, and the agent’s overall sentiment is represented by the average score across the five dimensions. This granular breakdown allows us to capture how different aspects of belief influence investment decisions and market outcomes over time.
Our results from the TwinMarket environment demonstrate that initial optimistic expectations often lead to increased buying activity, which in turn pushes stock prices higher. This reinforces investors’ confidence, creating a self-reinforcing feedback loop that drives prices further away from fundamental values. Eventually, as prices become unsustainable and deviate from fundamental support, the market corrects sharply, leading to a collapse, as shown in Figure 4. This observed pattern closely aligns with the classic bubble dynamics described in financial literature [31], highlighting the role of collective belief formation in driving both speculative booms and market crashes.
4.3 Macro-analysis: Stylized Facts
Figure 5: (a) Fat-tailed return distributions, (b) the leverage effect, (c) the volume-return relationship, and (d) volatility clustering. The results demonstrate that TwinMarket successfully replicates real market phenomena.
Existing studies have encoded key behavioral traits of market participants—such as the interplay between fundamentalist and chartist trading, herding tendencies, and decision-making heterogeneity—into simplified rules, such as probability-based strategy switching and structural stochastic volatility (SSV) [32, 33, 34]. This abstraction has been shown to effectively reproduce stylized facts of financial markets, including fat-tailed return distributions and volatility clustering [35, 36, 37].
LLM-based users inherently integrate the aforementioned behavioral traits, as LLMs have been demonstrated to analyze financial markets using both fundamental and technical metrics [38, 39]. These agents naturally exhibit herding tendencies [21]. To further capture the heterogeneity of market participants, we construct personalized user profiles for each users. These profiles enable agents to autonomously exhibit both rational investment strategies and irrational behavioral biases during decision-making, mirroring real-world investors. Further details on these dynamics are provided in Appendix E.1.
We compared the key stylized facts [^3] Stylized facts are empirical statistical patterns observed in real-world financial markets. They are commonly used to evaluate the effectiveness of simulated markets in replicating real-world dynamics[35]. of TwinMarket with historical market data, ensuring that our simulations accurately reflect real-world market behaviors. Key comparisons of these stylized facts are shown in Figure 5.
  • Non-Normality [40]: The log returns exhibit a sharp peak and fat tails.
  • Leverage Effect [41]: Negative returns show weak autocorrelation, indicating persistence in downturns.
  • Volume-Return Relationship [42]: Trading volume is positively correlated with returns, which typically reflects collective behavior among investors.
  • Volatility Clustering [43]: GARCH analysis [^4] GARCH model is widely used to model time-varying volatility in financial markets [44]. In the GARCH framework, $\alpha$ represents the sensitivity of current volatility to past returns or shocks, while $\beta$ captures the persistence of past volatility. shows that $\alpha+\beta$ is close to 1, indicating high persistence of volatility shocks.
The results show that TwinMarket, through a refined modeling of market participants and the market environment without relying on any rule-based techniques, is able to capture the macro-level dynamics of financial markets.
5 Emergence of Group Behaviors
Figure 6: Impact of rumor propagation on user behavior and market dynamics. (a) Users’ belief scores significantly decline under rumor exposure. (b) Users exhibit a stronger tendency to sell, leading to a sharp increase in the Sell/Buy ratio. (c) The collective reaction to rumors triggers a market downturn, causing a notable drop in stock prices.
In financial markets, individual decisions are not made in isolation but are embedded within a complex social network. Price feedback, social interactions, and information flow create a dynamic feedback loop that amplifies or dampens individual actions. Through mechanisms like social influence and herd behavior, individual choices are continually reshaped by the collective actions and expectations of others, leading to the emergence of market trends and fluctuations. Through detailed modeling of individual agent behavior and their social interactions, our simulation demonstrates how macroscopic market phenomena emerge organically, validating core tenets of economic theory and mirroring empirical market observations, without imposing artificial constraints or pre-defined rules.
5.1 Information Propagation
Previous financial research has explored various phenomena, such as echo effects [45, 46], that arise from information propagation. To further investigate the mechanisms behind this propagation and its potential impact on the market, we begin by selecting the most important news or reports each day, which are then forcibly pushed to high-centrality users. This approach reflects the information asymmetry present in the market environment and enables us to study the dynamics of information flow.
In addition, to compare and validate the impact of information on the market, we also replace a set of news with rumors, which are shown in Appendix D.4. These rumors are more extreme and sensational compared to regular news, potentially amplifying the effects of information flow in the market. Further explanations and results will bee illustrated in Appendix D.2.
Figure 7: The adjacency matrix heatmap highlights distinct density patterns, particularly in the red-boxed regions, indicating strong behavioral polarization among user groups.
Intimation & Polarization.
@@ 5.1 Information Propagation
Figure 8: Average likes trends in high-centrality users and low-centrality users.
@@
In TwinMarket, graph edges represent behavioral similarity, meaning users closely connected exhibit stronger intimation effects. As shown in Figure 8, high-degree users receive more upvotes, indicating their greater influence within the network. These opinion leaders [47] shape decision-making, fostering homogeneity and leading to polarization, as depicted in Figure 7. The network’s adjacency matrix highlights clusters of trading similarity, further intensified under rumor conditions. This suggests that biased or extreme information spreads more rapidly in tightly connected groups, reinforcing polarization [48, 49].
Belief Divergence.
Figure 6(a) shows that beliefs in normal and rumor scenarios increasingly diverge [^5] In rumor-exposed scenarios, agents maintained 27.5% lower valuations than the control group., forming distinct echo chambers [50, 51]. This occurs as negative rumors replace normal news, shifting users’ market expectations [52]. Rumor-exposed users become more pessimistic, reassessing asset values downward, reinforcing their isolation, and further distancing themselves from unaffected groups[53]. This process exemplifies how biased information amplifies belief divergence and strengthens echo chambers in financial markets.
Trading Volatility.
Trading behavior in TwinMarket is primarily driven by users’ intentions and beliefs about the market, which in turn influence their buying and selling decisions. When belief evaluations fluctuate, particularly in response to rumors, these changes significantly impact trading activity [54, 55]. As shown in Figure 6(b), under the influence of negative rumors, the proportion of sell orders increased significantly. The Sell/Buy ratio to surge to 2.02$\times$ the baseline6660.997 vs 0.495, indicating amplified panic-driven selling behavior.
Market Turbulence.
The compounding effect of shifting beliefs and rumor-driven uncertainty leads to significant market instability. Figure 6(c) shows that while normal market conditions remained stable, rumor-exposed markets suffered a sharp decline. This drop reflects panic-driven sell-offs and overreaction to misinformation [56]. As negative sentiment spreads, confidence erodes, reinforcing self-perpetuating downturns and destabilizing asset values.
5.2 Scaling Up Design
Figure 9: Simulated and real index price comparison.
To demonstrate the scalability of our framework, we expanded our simulation to include 1,000 agents, a scale typically challenging for traditional agent-based models. The results, as depicted in Figure 9, reveal a very similar overall trend and movement throughout the five-month simulation period. This finding aligns with research showing that complex systems can exhibit similar macroscopic behavior when scaled up [57], even if the individual interactions at the micro-level become more numerous. In our case, it appears that the overall market dynamics can be well reproduced through emergent behaviors of LLM-powered agents. A more comprehensive quantitative evaluation, including metrics such as RMSE, MAE, and lag analysis, are detailed in Appendix E.2.
@@ Appendix A Related Work
Large language models (LLMs) excel at simulating complex human behaviors, from rational decision-making and financial market analysis to replicating behavioral biases like herding and overconfidence [3, 39, 22, 21]. Beyond behavioral simulation, they enable fine-grained modeling of individual actions and large-scale emergent phenomena [1, 2]. LLMs naturally exhibit demographic diversity, such as gender, age, and education, while engaging in rich social interactions [22, 66]. By automating text analysis, generating experimental stimuli, and designing multi-agent systems, they offer powerful tools for social science research [67, 68, 4].
Appendix B TwinMarket Statistics
B.1 SSE 50 and 10 Aggregated Index

### v3

@@ 5.1 Micro-level Emergence: Self-fulfilling Prophecy
A self-fulfilling prophecy describes a process where individual expectations collectively influence behaviors in ways that ultimately validate the original expectations [40]. This mechanism exemplifies micro-level emergence: localized belief updates and actions aggregate into systemic market outcomes. In our BDI framework, agents continuously revise beliefs, which guide their trading decisions and shape market dynamics. TwinMarket simulations reveal that early optimism amplifies buying pressure, driving prices upward and reinforcing collective confidence. This positive feedback loop detaches prices from fundamentals until a sharp correction ensues. As shown in Figure 10, belief and price co-evolve, replicating the speculative boom-bust cycle [41] and highlighting how collective belief formation drives macro-level patterns.
5.2 Macro-level Emergence: Information Propagation
@@
Figure 11: Average likes trends among different centrality users.
Previous work highlights how information propagation produces market-wide echo effects [42, 43]. To probe the same mechanism, we identify the day’s most salient news items and deliver them directly to users with the highest network centrality, reproducing the information asymmetry characteristic of real markets. We then rerun the experiment with a parallel set of exaggerated, rumor-like headlines (We list some examples in Appendix G.3.) to stress-test the system. Observing how these targeted, micro-level interventions cascade through the network and reshape aggregate prices and volumes allows us to treat information diffusion as a source of macro-level emergence: the collective dynamics that arise cannot be inferred by examining any single agent in isolation, but only by tracking the system-wide response to the injected signals.
@@
Intimation & Polarization. In TwinMarket, graph edges represent behavioral similarity, meaning users closely connected exhibit stronger intimation effects. As shown in Figure 11, high-degree users receive more upvotes, indicating their greater influence within the network. These opinion leaders [44] shape decision-making, fostering homogeneity and leading to polarization, as depicted in Figure 12 (further explored in Appendix D.4). The network’s adjacency matrix highlights clusters of trading similarity, further intensified under rumor conditions. This suggests that biased or extreme information spreads more rapidly in tightly connected groups, reinforcing polarization [45, 46].
Figure 12: Heatmap of the user trading similarity adjacency matrix. Higher density indicates greater behavioral similarity.
Belief Divergence. Figure 13(a) illustrates how belief trajectories diverge over time between the control and rumor-exposed groups. [^4] On average, rumor-exposed agents assigned asset valuations 27.5% lower than those in the baseline. Replacing neutral news with negative rumors pushes expectations downward and reinforces echo chamber dynamics [47, 48]. Rumor-affected users become increasingly pessimistic, interact primarily with like-minded peers, and distance themselves from the broader population [49, 50]. This process intensifies belief heterogeneity and entrenches informational silos in the market.
Trading Volatility. Rumors do not merely shift beliefs, they also alter market behavior. As valuations decline, agents transition from buying to selling, triggering sharp changes in order flow [51, 52]. Figure 13(b) shows that the sell-to-buy ratio nearly doubles under rumor exposure, rising from 0.495 to 0.997. This dramatic increase reflects panic-driven responses to uncertainty and exemplifies how misinformation can amplify volatility in financial systems.
Figure 13: Rumor propagation impacts user beliefs and market trends. (a) Users’ belief scores decline under rumor. (b) Users exhibit a stronger tendency to sell, and Sell/Buy ratio increase. (c) The collective reaction to rumors triggers a market downturn, causing stock prices drop-down.
Market Turbulence. The cumulative effect of belief erosion and reactive trading destabilizes the broader market. Figure 13(c) shows that while prices remain stable in the control setting, rumor-exposed markets experience a pronounced decline. This reflects overreaction and herding behavior under uncertainty [53]. As negative sentiment spreads, confidence collapses, triggering self-reinforcing downturns that destabilize asset values and contribute to systemic fragility.
6 Ablation Study
To assess the contributions of TwinMarket’s core design elements, we conduct an ablation study comparing the full model against two variants: (1) w/o BDI, which removes the belief-desire-intention module, making agents purely reactive; and (2) w/o Hetero., which eliminates agent heterogeneity by assigning uniform strategies and biases. As shown in Table 5, removing either component degrades performance across predictive accuracy (RMSE, MAE, correlation) and financial realism (kurtosis, leverage effect, GARCH parameters).
Table 5: Ablation study on the effects of the BDI framework and agent heterogeneity. $\uparrow$: higher is better, $\downarrow$: lower is better.
Model Variant RMSE $\downarrow$ MAE $\downarrow$ Corr. $\uparrow$ Kurtosis $\uparrow$ Lev. Effect $\uparrow$ GARCH $\alpha$uparrow GARCH $\beta$uparrow
TwinMarket 0.02 0.02 0.77 5.24 0.11 0.06 0.83
TwinMarket w/o BDI 0.07 0.05 0.34 4.25 0.05 0.03 0.76
TwinMarket w/o Hetero. 0.09 0.08 -0.61 3.58 0.13 0.05 0.85
The BDI ablation notably weakens the model’s ability to generate stylized facts, such as fat tails and asymmetric volatility, highlighting the importance of cognitive reasoning in agent behavior. In contrast, removing heterogeneity leads to a sharper collapse in realism, with near-Gaussian returns and negative correlation, underscoring the necessity of strategic diversity for emergent market complexity. Together, these results confirm that both BDI framework and agent heterogeneity are essential for replicating real-world market dynamics. Additional ablation studies appear in Appendix D.2.
7 Scalability
[…]
Figure 15: Simulation result of 1,000 agents.
To assess the potential of data-driven scaling as a performance enhancement strategy, we investigate how varying levels of agent activation affect TwinMarket’s predictive performance. We simulate markets with 10%, 20%, 40%, and 80% of agents actively trading each day, [^5] Total 1,000 agents remain socially interactive; only a subset engage in trading. The experiments are conducted from June 15 to July 15, 2023. while keeping other conditions fixed. As shown in Figure 15, higher activation levels lead to increased trading volume and stronger collective dynamics. Figure 15 further shows that TwinMarket, even with 1,000 agents, can reproduce realistic market trends over extended periods. Together, these results reveal a clear scaling law: greater activation proportion enhances both behavioral realism and market-level alignment. This highlights TwinMarket’s ability to capture macro-level market dynamics as emergent outcomes of micro-level agent interactions.
@@ Appendix A Related Work
Large language models (LLMs) excel at simulating complex human behaviors, from rational decision-making and financial market analysis to replicating behavioral biases like herding and overconfidence [3, 63, 23, 22]. Beyond behavioral simulation, they enable fine-grained modeling of individual actions and large-scale emergent phenomena [1, 2]. For instance, LLM-based multi-agent systems like TrendSim have been developed to simulate specific social media dynamics such as the spread of trending topics and the impact of poisoning attacks, incorporating features like time-aware interactions and human-like agent profiles with memory and perception modules [64]. The capacity for large-scale simulation is further demonstrated by platforms such as AgentSociety, which employs LLM-driven agents to investigate complex social issues like polarization and the effects of universal basic income by modeling agent minds with emotions, needs, and cognition [65]. Similarly, SocioVerse aims to create world models for social simulation by leveraging pools of millions of real-world user profiles to align simulations in politics, news, and economics with reality [66]. General-purpose platforms like GenSim are also emerging, supporting simulations with multi-agents and incorporating error-correction mechanisms to enhance the robustness and adaptability of customized social scenarios [67]. LLMs naturally exhibit demographic diversity, such as gender, age, and education, while engaging in rich social interactions [23, 68]. By automating text analysis, generating experimental stimuli, and designing multi-agent systems, they offer powerful tools for social science research [4, 69, 70]. Crucially, the reliability and consistency of these LLM-based simulations are active areas of research, leading to the development of specialized datasets and methods to systematically evaluate and improve how LLMs embody simulated roles and maintain persona fidelity [71]. The application of such LLM-powered agents to financial market simulation, as explored in our work, promises to capture more nuanced trader behaviors and market dynamics.
Appendix B Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix C Limitation
This research is conducted solely for academic purposes. The simulation results and models presented are not intended for real-world trading or financial decision-making.
Appendix D More Experimental Analysis
D.1 Social Network Sensitive Analysis
To assess the robustness of our social network construction, we conduct a sensitivity analysis by varying two key parameters: the similarity threshold and the time decay factor. All tests are performed on a social network with 100 users.
We first vary the similarity threshold while fixing the time decay factor at 0.5. A lower threshold allows weakly aligned users to connect, resulting in denser graphs with higher noise. A higher threshold filters out weaker ties, making the graph sparser but more locally cohesive. We then vary the time decay factor $\lambda$ while fixing the similarity threshold at 0.2. Smaller values preserve long-term behavioral influence, while larger values emphasize recent activity. As shown in Figure 16, higher parameter values generally lead to sparser, less modular, and less clustered graphs. The size of the largest connected component remains stable across different settings. We find that using a similarity threshold of 0.2 and a decay factor of 0.5 maintains moderate connectivity and avoids over-saturation, supporting meaningful information propagation.
Figure 16: Sensitivity analysis of the graph structure under different similarity thresholds (top) and time decay factors (bottom). Higher parameter values generally lead to sparser, less modular, and less clustered graphs. The largest connected component remains stable across settings.
D.2 More Ablation Study
To assess the robustness of our TwinMarket framework and understand the impact of various components and parameters, we conducted several other ablation studies. This section details these investigations.
D.2.1 Impact of LLM Temperature on Agent Behavior
Table 6: Ablation on LLM Temperature
Temp. Sells Buys Avg. Volume
0.3 2369 3581 4818.31
0.7 2402 3583 4794.72
1.0 2403 3593 5580.91
1.3 (our setting) 2374 3586 6071.78
The temperature parameter in LLMs controls the randomness and creativity of the generated text. A lower temperature results in more deterministic and focused outputs, while a higher temperature leads to more diverse and potentially surprising responses. To evaluate how this parameter influences agent behavior and market dynamics within TwinMarket, we ran simulations with varying temperature settings for the LLM powering our agents.
The results are summarized in Table 6. It is shown that core trading decisions (total “Sells” and “Buys”) remain remarkably stable across various temperature settings while the “Avg. Volume” per transaction generally increases with rising temperatures. This suggests that while the decision to trade remains stable, higher temperatures, which encourage more diverse LLM reasoning, might lead to agents expressing decisions with greater perceived conviction or scale, thus increasing average transaction sizes. Despite LLM generation parameters affecting textual outputs, our simulation framework demonstrates stable core decision-making under reasonable parameter variations. This validates that the observed emergent behavior stems from spontaneous dynamics and the framework’s superiority, without posing information leakage risks.
D.2.2 Sector-Specific Dynamics: Simulation Across Different Industries
To capture the heterogeneous nature of real-world financial markets, TwinMarket simulates market dynamics across 10 aggregated industry indices (detailed in Appendix E.1.1). This ablation study assesses whether our framework generates distinct and plausible dynamics for different sectors. We analyzed key statistical properties of their simulated returns, focusing on characteristics indicative of real-world financial time series, such as fat tails (kurtosis) and volatility clustering (GARCH parameters $\alpha$ and $\beta$).
Table 7: Ablation on Industries
Industry (Index) Kurtosis GARCH-$\alpha$ GARCH-$\beta$
SSE 50 5.24 0.06 0.83
Consumer Goods Index (CGEI) 3.06 0.05 0.71
Technology and Telecommunications Index (TTEI) 6.53 0.09 0.96
The results are summarized in Table 7. The simulated differences – a more volatile and fat-tailed tech sector (TTEI) versus a stabler consumer goods sector (CGEI) – align with common real-world observations and expectations. This correspondence demonstrates TwinMarket’s capability to generate differentiated and plausible dynamics across various simulated industries, supporting the framework’s applicability for studying multi-asset markets where sector-specific behaviors are important.
D.2.3 Impact of Social Interaction on Agent Decision-Making
To assess the influence of real-time social interactions versus historical data on agent decision-making, we We compared the performance of the full TwinMarket model with a variant where agent social interactions (the social meida platform) were disabled. Both simulations ran from 2023-06-15 to 2023-08-15 with 100 agents.
Table 8: Ablation on Social Interaction
Model RMSE MAE Correlation
TwinMarket 0.02 0.02 0.77
TwinMarket w/o Interaction 0.18 0.07 -0.50
The results, presented in Table 8, demonstrate a significant degradation in performance when social interactions are removed. This suggests that agents in our simulation prioritize and are heavily influenced by real-time social interactions and emergent group dynamics rather than solely relying on historical data. This finding also indicates that potential data leakage from historical news or announcements has a negligible impact on our primary results, as the dynamic social component plays a more dominant role in shaping agent decisions and overall market outcomes.
D.3 Reproducibility and Stability Validation
To evaluate the reproducibility of emergent phenomena and the stability of the simulation framework under varied conditions, we performed multiple simulation runs (3 iterations each) for a consistent period from June 15th to August 15th, 2023, with two different models as the backbone for agent modeling.
Table 9: Reproducibility and Stability: Comparison of Aggregated Simulation Metrics Across Different Foundation Models. Data are reported as mean $\pm$ relative standard deviation (RSD), with RSDs shown as percentages (i.e., mean $\pm$ RSD%).
Backbone Model RMSE MAE Kurtosis Lev. Effect GARCH-$\alpha$ GARCH-$\beta$
GPT-4o (3 runs) $0.023\pm 31.13\%$ $0.026\pm 36.21\%$ $4.433\pm 20.85\%$ $0.060\pm 43.01\%$ $0.047\pm 63.17\%$ $0.780\pm 27.41\%$
Gemini-1.5-Flash (3 runs) $0.024\pm 16.34\%$ $0.020\pm 23.12\%$ $4.886\pm 8.25\%$ $0.108\pm 16.52\%$ $0.032\pm 61.67\%$ $0.897\pm 13.26\%$
TwinMarket (1 run) $0.020$ $0.020$ $5.24$ $0.11$ $0.06$ $0.83$
Table 9 presents a comparison of key performance and market characteristic metrics across the multiple runs with different LLM backbones, against our established TwinMarket framework. The quantitative analysis demonstrates that TwinMarket simulations, irrespective of the specific advanced LLM used, reliably capture key aspects of real-world market dynamics. The low error metrics (RMSE & MAE) indicate a strong correspondence between simulated and historical price movements.
Notably, our experiments consistently demonstrate that crucial market phenomena, such as fat-tailed return distributions (indicated by Kurtosis values significantly greater than 3), volatility clustering (evidenced by the GARCH $\alpha$ and $\beta$ parameters where $\alpha>0$, $\beta>0$, and $\alpha+\beta$ is close to 1), and the leverage effect – emerge reliably across all simulation runs and with both LLM backbones.
While the specific parameter values for these stylized facts exhibit some variability across runs and between models, the qualitative presence and general characteristics of these stylized facts remain stable. This observation is consistent with real financial markets, where core statistical patterns persist despite short-term fluctuations or differences in the precise estimation of individual metrics. The observed variability in parameters like GARCH coefficients is expected, reflecting the inherent stochasticity present in complex financial systems and in the LLM’s generative process. As highlighted in prior literature [72, 73], the primary validation focuses on whether the model can consistently reproduce the qualitative stylized facts and emergent behaviors across simulations, rather than demanding numerical stability of every single estimated parameter to an exact value. Our results affirm this capability for TwinMarket.
D.4 Emergence of Opinion Leaders
This section provides additional validation for TwinMarket by demonstrating its capability to simulate the emergence of opinion leaders, a key phenomenon in social networks[74]. This emergence is driven by the platform’s information aggregation mechanism and the dynamics of social interaction, where engaging content gains visibility and influences user behavior.
Within TwinMarket, posts that resonate strongly with the broader community achieve wider dissemination. As users interact with content and potentially imitate the behaviors or adopt the perspectives of influential posters, their social connections can become more centralized around these key individuals. This self-reinforcing process naturally leads to the emergence of opinion leaders—users whose posts and opinions exert a disproportionately larger impact on the network.
(a) Post-propagation chains in the social network. Nodes with high connectivity and wide reach emerge as potential opinion leaders.
(b) Trend of repost counts in the 100-user network. Some users show much greater reach, signaling emergent influence.
Figure 17: Illustration of information diffusion in TwinMarket. (a) Visualizes post-propagation paths; (b) shows how repost counts evolve over time, highlighting key opinion leaders.
As illustrated in Figure 17(a), users who achieve high-degree centrality effectively gain influence as their posts receive increased engagement, leading to broader information diffusion. This effect is further quantified in Figure 17(b), which shows that a small subset of users experiences significantly higher repost counts over time. This skewed distribution of engagement reinforces their role as key opinion leaders. Their ability to shape market sentiment and drive collective decision-making underscores the impact of social influence in financial markets, where information cascades can significantly alter trading behaviors and reshape network structures, which aligns with observed real-world social dynamics and further validates its design.
D.5 Qualitative Validation of Micro-Level Behaviors
To validate the realism of individual agent behaviors within TwinMarket, this section presents qualitative examples of agent decision-making logs. These examples, generated through interactions guided by their initialized personas and BDI-driven cognitive process, demonstrate the agents’ capacity to autonomously exhibit a range of behaviors observed in real-world investors.
D.5.1 Demonstration of Investment Analysis Approaches
TwinMarket agents are capable of employing recognized investment analysis techniques, as discussed in Appendix F.1. As shown in Figure 18, an agent demonstrates Fundamental Analysis by considering valuation metrics, long-term outlook, and national strategy to justify increasing its position. Similarly, Figure 19 illustrates an agent engaging in Technical Analysis, focusing on volume-price action, moving averages, and momentum to make trading decisions, while also considering stop-loss levels for risk management.
Figure 18: Example of Fundamental Analysis.
Figure 19: Example of Technical Analysis.
D.5.2 Exhibition of Behavioral Biases
Beyond structured analytical approaches, TwinMarket agents also manifest common behavioral tendencies and socially influenced actions. The Disposition Effect is illustrated in Figure 20, with an agent intending to sell a winning position prematurely. Figure 21 showcases Lottery Preference, where an agent favors a high-risk, speculative investment despite its poor past performance. Underdiversification is evident in Figure 22, as an agent admits to concentrating on familiar assets while neglecting others.
Figure 23 provides an example of Overconfidence, where an agent displays excessive conviction in its trading strategy. Furthermore, Figure 24 captures an instance of Herd Behavior, where an agent’s trading decision is swayed by perceived widespread forum sentiment and low market volume, overriding its typical technical approach.
Figure 20: Examples of Disposition Effect.
Figure 21: Examples of Lottery Preference.
Figure 22: Examples of Underdiversification.
Figure 23: Examples of Overconfidencce.
Figure 24: Examples of Herd Behavior.
Appendix E Technical Details I: Data Statistics
E.1 Stock Data Details
E.1.1 SSE 50 and 10 Aggregated Index